A calorimeter contains 250 gram of water at odegree c and 15 gm of ice at o degree c now steam at 100 degree cis 5 into it continuously till the temperature of the content become 15 degrees calculate the amount of stamp required if water equivalent of the calorimeter loss suffered due to radiation is negligible

Grade:12th pass

1 Answers

Shiv Sagar
4 years ago
Since final temperature is 15 degrees, all the substances reach that state.
 
Heat absorbed by 250g water at 0 degrees = 250 x 1 x (15 – 0)  = 3750 cal
Heat absorbed by 15g ice = 15 x 80 [latent heat] + 15 x 1 x (15 – 0) = 1425 cal
 
This heat will be provided by an amout “m” of steam that we poured and have to calculate.
 
Heat lost by “m” mass of steam = m x 540 + m x 1 x (100 – 15) = 615 x m
This is equal to 3750 + 1425
3750 +1425 = 615m,
m = 8.414g